# o m ∥#2 ∥_#1

### v1

@@ Appendix E Hyperparameters
For our neural network experiments, we use the implementation of PQN available via CleanRL [HDYB+22]. We report the hyperparameters that we used in Table 1. We found that from the original implementation we need to make minor modifications that do not lead to decrease in performance of the baseline to get competitive performance with quantization. Precisely, we change the default exploration rate from $0.1$ to $0.3$ because quantization leads to lower policy churn which has been shown to increase exploration [SBQO22]. This is an intended side effect. While one might be tempted to argue that exploration from churn is good, it is uncontrolled as we do not know how our networks change. Instead, quantization leads to lower churn and the modeler can control the rate of exploration directly. Furthermore, we change the optimizer to use decoupled weight decay [LH19].

### v2

@@ Appendix E Hyperparameters
For our neural network experiments, we use the implementation of PQN available via CleanRL [HDYB+22]. We report the hyperparameters that we used in Table 1. We found that from the original implementation we need to make minor modifications that do not lead to decrease in performance of the baseline to get competitive performance with quantization. Precisely, we change the default exploration rate from $0.1$ to $0.3$ because quantization leads to lower policy churn which has been shown to increase exploration [SBQO22]. This is an intended side effect. While one might be tempted to argue that exploration from churn is good, it is uncontrolled as we do not know how our networks change. Instead, quantization leads to lower churn and the modeler can control the rate of exploration directly. Furthermore, we change the optimizer to use decoupled weight decay [LH19]. All outputs are rounded to a multiple of $0.4$ during quantization.